OpenStudy (anonymous):

Hey guys can I get some help?! I need step by step instruction when evaluating sin(Cos^-1(4/5)+Tan^1(3/4)). What is an identity used in the solution? a. sin ( x + y) = sin x cos y - cos x sin y b. sin ( x + y) = cos x cos y + sin x sin y c. sin ( x + y) = sin x cos y + cos x sin y d. sin ( x + y) = cos x cos y - sin x sin y

hero (hero):

\(\cos^{-1}(\frac{4}{5}) + \tan^{-1}(\frac{3}{4})\) Is that right?

OpenStudy (anonymous):

Yes, that's right

OpenStudy (anonymous):

@Hero

hero (hero):

I'm stumped.

OpenStudy (anonymous):

sin(Cos^-1(4/5)+Tan^1(3/4)) I think you forgot the sin in front

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

Identity that will be used in this solution will be the first one i.e. A

OpenStudy (anonymous):

sorry my Bad you have to go for third Identity i.e. C

OpenStudy (anonymous):

Oh, can you please show me how you got that??

OpenStudy (anonymous):

you know that formula for Identity Sin(A+B) = CosA*SinB+SinA*CosB

OpenStudy (anonymous):

Can i give an idea here if you guys don't mind

OpenStudy (anonymous):

then we have to use the same formula @godorovg go ahead

OpenStudy (anonymous):

Oh O.K. Yes please do @godorovg

OpenStudy (anonymous):

we know sin/cos = tan is right/ so why want sub in and slove in this manner

OpenStudy (anonymous):

and work backwards just a thought

OpenStudy (anonymous):

like this tan a-b/ cos a-b =sin a-b /2/sin a-b2 =sin a-b /ctn C/2 and so on

OpenStudy (anonymous):

thoughts here

OpenStudy (anonymous):

Oh ok. And that gives us c. as well?

OpenStudy (anonymous):

@godorovg

OpenStudy (anonymous):

yes it should if done right

OpenStudy (anonymous):

Thank you for your input!

OpenStudy (anonymous):

does that make sense to you??